# RemoteReasoner: Towards Unifying Geospatial Reasoning Workflow

### v2

@@ Abstract
Remote sensing imagery presents vast, inherently unstructured spatial data, necessitating sophisticated reasoning to interpret complex user intents and contextual relationships beyond simple recognition tasks. In this paper, we aim to construct an Earth observation workflow to handle complex queries by reasoning about spatial context and user intent. As a reasoning workflow, it should autonomously explore and construct its own inference paths, rather than being confined to predefined ground-truth sequences. Ideally, its architecture ought to be unified yet generalized, possessing capabilities to perform diverse reasoning tasks through one model without requiring additional fine-tuning. Existing remote sensing approaches rely on supervised fine-tuning paradigms and task-specific heads, limiting both autonomous reasoning and unified generalization. To this end, we propose RemoteReasoner, a unified workflow for geospatial reasoning. The design of RemoteReasoner integrates a multi-modal large language model (MLLM) for interpreting user instructions and localizing targets, together with task transformation strategies that enable multi-granularity tasks, including object-, region-, and pixel-level. In contrast to existing methods, our framework is trained with reinforcement learning (RL) to endow the MLLM sufficient reasoning autonomy. At the inference stage, our transformation strategies enable diverse task output formats without requiring task-specific decoders or further fine-tuning. Experiments demonstrated that RemoteReasoner achieves state-of-the-art (SOTA) performance across multi-granularity reasoning tasks. Furthermore, it retains the MLLM’s inherent generalization capability, demonstrating robust performance on unseen tasks and out-of-distribution categories.
Keywords: Remote sensing, Reasoning workflow, Reinforcement learning
= Projects: Work in progress.
= Contact: yaoliang@hhu.edu.cn
@@ 1 Introduction
Remote sensing analysis has evolved beyond simple object detection Zou et al. (2023), Liu et al. (2025a) or classification Adegun et al. (2023), increasingly requiring stronger capabilities to support complex real-world decision-making. Although existing approaches Yao et al. (2025b), Liu et al. (2024a) are effective for multiple tasks, they struggle to interpret nuanced user intent and infer implicit spatial relationships for intricate queries. Reasoning Jaech et al. (2024), Xu et al. (2025) bridges this gap by enabling systems to dynamically decipher the underlying goals of ambiguous or high-level instructions (e.g., “Find areas at risk of flooding near critical infrastructure after heavy rain"). Such a geospatial workflow would conveniently accommodate varied geoscience applications Zhao et al. (2024).
@@
Figure 1: Comparison between Existing Frameworks and RemoteReasoner. Existing remote sensing reasoning approach (e.g., , SegEarth-R1) requires SFT with annotated reasoning processes, and is limited to single-task outputs and task decoder. In contrast, our framework supports unsupervised reasoning and multi-granularity tasks.
@@
Numerous efforts Lai et al. (2024), Ren et al. (2024), Chen et al. (2024b) have explored reasoning tasks in the natural image domain. For instance, LISA Lai et al. (2024) utilizes multi-modal large language models (MLLMs) Carolan et al. (2024) for reasoning-driven segmentation tasks. However, these approaches rely on supervised fine-tuning (SFT) paradigms to train the models. Such frameworks constrain model autonomy, ultimately compromising generalization capabilities. This situation can be manifested in substantial performance degradation on out-of-distribution samples and catastrophic forgetting of general capabilities Liu et al. (2025b). Similarly, the remote sensing community has seen pioneering works Li et al. (2025), Zhang et al. (2023), Wang et al. (2024). SegEarth-R1 Li et al. (2025) introduces a geospatial pixel reasoning task that integrates a hierarchical MLLM and a mask decoder to generate both reasoning processes and segmentation outputs. However, it remains constrained by SFT paradigms and exclusively supports segmentation tasks.
@@
These limitations motivate us to design a reasoning workflow that autonomously interprets user intent through self-directed thinking. It should also efficiently generate geospatial interpretations in diverse output formats, which are tailored for various downstream intelligence analysis. Firstly, the requirement for self-thinking should be fulfilled through an open-ended learning paradigm. This aligns naturally with the emerging reinforcement learning (RL) frameworks Guo et al. (2025), Shen et al. (2025). Rather than SFT approaches that demand strictly annotated reasoning processes, RL explores the model’s self-evolution potential to cultivate reasoning capabilities. Secondly, to flexibly support multi-granularity tasks, we should avoid redundant task decoders and training through decision-level task transformation.
@@
Followed by these insights, we propose RemoteReasoner, a reasoning workflow that integrates a robust MLLM and a flexible inference strategy. As shown in Fig. 1, building upon the pixel-level reasoning task proposed in SegEarth-R1 Li et al. (2025), RemoteReasoner further implements two novel tasks through autonomous reasoning: 1) region-level: geospatial region reasoning, and 2) object-level: geospatial contour reasoning. These three distinct tasks can be achieved through a single MLLM inference followed by a unified decision-level task transformation pipeline, significantly improving output efficiency for multi-granularity reasoning tasks.
@@
During training, we utilize Group Relative Policy Optimization (GRPO) Shao et al. (2024) to fine-tune the reasoning model. Benefiting from our sophisticated reward function that balances output formats, localization accuracy, and quantity, RemoteReasoner demonstrates strong capabilities in autonomous reasoning and precise region-level localization. This training approach maintains the inherent generalization capability of MLLM, thereby endowing it with the ability to recognize out-of-distribution categories. During inference, we design a task transformation pipeline inspired by RemoteSAM Yao et al. (2025b), which enables the model to perform object- and pixel-level reasoning tasks, eliminating redundant computations across granularities.
@@
Extensive experimental results demonstrate that RemoteReasoner achieves state-of-the-art performance on multi-granularity reasoning tasks, evidencing its enhanced autonomous reasoning capabilities and superior generalization. For instance, RemoteReasoner achieves 3.67% accuracy improvement over SegEarth-R1 in pixel-level geospatial reasoning tasks. On unseen tasks (e.g., , Referring Segmentation), our framework significantly outperforms competing models (e.g., , PixelLM, LISA) in performance.
@@ 2.1 Reasoning in Large Language Models
Recent efforts Jaech et al. (2024), Guo et al. (2025), Ren et al. (2024), Lai et al. (2024), Li et al. (2025) have shown that reasoning is central to aligning large language models (LLMs) with complex instructions. Models like OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) incorporate chain-of-thought(CoT) Wei et al. (2022) reasoning into their generation process, enabling flexible and interpretable multi-step inference. Both adopt reinforcement learning (RL) Kaelbling et al. (1996), Li (2017), Arulkumaran et al. (2017) to induce generalizable reasoning behaviors—OpenAI-o1 through RLHF and CoT monitoring, and DeepSeek-R1 via large-scale RL from scratch followed by multi-stage pretraining. These approaches exhibit strong reasoning capabilities but remain limited to natural language and vision tasks.
In remote sensing, models such as PixelLM Ren et al. (2024), LISA Lai et al. (2024), and SegEarth-R1 Li et al. (2025) extend reasoning to spatial domains. While effective for pixel- or region-level inference, they often support narrow task types and fixed granularity. Notably, SegEarth-R1 relies on supervised fine-tuning, making it less robust to distribution shifts and less adaptable in open-ended settings.
These limitations underscore the need for a more versatile geospatial framework. To this end, we propose RemoteReasoner, which supports unified and efficient multi-task, multi-granularity remote sensing understanding.
@@ 2.2 Remote Sensing Multi-modal Models
Recent remote sensing multimodal models Muhtar et al. (2024), Yao et al. (2025a), Zhan et al. (2024), Kuckreja et al. (2024), Yao et al. (2025b) demonstrate strong capabilities in handling diverse tasks across spatial granularities. Falcon Yao et al. (2025a) supports 14 vision-language tasks via multi-resolution encoding; GeoChat Kuckreja et al. (2024) and LHRS-Bot Muhtar et al. (2024) incorporate spatial prompts and geographic priors for improved region-level grounding; RemoteSAM Yao et al. (2025b) adopts a unified encoder for dense pixel-level segmentation; SkyEyeGPT Zhan et al. (2024) reformulates task input/output as natural language for broad task generalization. While effective in geospatial understanding, these models lack robust reasoning capabilities, limiting performance under ambiguous or compositional prompts.
While existing models demonstrate strong perception capabilities, they lack the reasoning ability to interpret complex prompts or adapt across granularities. RemoteReasoner addresses this gap by incorporating autonomous reasoning, enabling precise execution of complex natural language instructions and strong generalization across multi-granularity geospatial tasks.
Figure 2: Overview of our RemoteReasoner. We utilize GPRO Shao et al. (2024) to explore model’s self-thinking capability. Then we design an inference workflow to perform multi-level reasoning tasks.
@@ 3.1 Data
The EarthReason Li et al. (2025) dataset is proposed alongside SegEarth-R1, containing 6 questions per sample with an average length of 20.86 words. It comprises 2,371 training, 1,135 validation, and 1,928 testing samples. However, this dataset is limited to pixel tasks. To address this, we generate detection and contour annotations through Mask2Bbox (M2B) Liu et al. (2024a) and Mask2Contour (M2C) Miao et al. (2025) strategies, respectively. Additionally, we design prompt templates combined with original questions and generated bounding boxes to form training samples. The segmentation and contour annotations are exclusively reserved for downstream testing evaluation.
@@ 3.2 Reasoning Model
We utilize Qwen2.5-VL-7B-instruct Bai et al. (2025) as our reasoning model $\mathcal{F}_{\theta}$. Despite Qwen2.5-VL’s superior performance in general-domain applications, direct adaptation to remote sensing imagery introduces domain knowledge discrepancy. Moreover, its reasoning capabilities require significant enhancement to address complex geospatial tasks. Therefore, we require fine-tuning it through reinforcement learning on remote sensing data.
@@ 3.3 Training
We utilize Group Relative Policy Optimization (GRPO) Shao et al. (2024) to train RemoteReasoner’s reasoning model. Unlike traditional reinforcement learning methods like PPO Schulman et al. (2017), GRPO directly compares groups of candidate outputs without an additional critic model. Given a query $Q$, GRPO samples N candidate outputs $o_{1},o_{2},...,o_{N}$ from the policy $\pi_{\theta}$. Then it computes each output $o_{i}$ through a Reward function $R(Q,o_{i})$. To determine the relative quality of the outputs, GRPO further normalizes the obtained rewards and subsequently derives the advantage:
@@
$\frac{1}{N}\sum_{i=1}^{N}\{min[c_{1}\cdot A_{i},c_{2}\cdot A_{i}]-\beta\mathbb{D}_{KL}[\pi_{\theta}||\pi_{ref}]\},$
@@ 3.4 Inference Workflow
After training, the reasoning model itself generates bounding boxes (bbox) indicating target regions. To further enhance the workflow and enable both pixel-level and object-level reasoning tasks, we need to leverage established tools. Specifically, we first input the bbox and its centroid coordinates to the SAM2 Ravi et al. (2024) to generate pixel-level masks. Subsequently, morphological operations are applied to analyze the mask images, extracting the outermost polygonal vertices to obtain target contours. The mathematical formulation of this process is presented as follows:
@@ 4 Experiments
LISA Lai et al. (2024) 57.39 59.10 61.04 60.88
PixelLM Ren et al. (2024) 57.79 59.22 57.94 60.01
PSALM Zhang et al. (2024) 62.03 64.61 66.61 68.30
SegEarth-R1 Li et al. (2025) 64.13 68.25 68.60 70.75
@@
DeepSeek-VL2-tiny Guo et al. (2025) 12.08 12.67 17.51 18.62
GeoChat Kuckreja et al. (2024) 10.10 8.89 12.57 11.44
Qwen2.5-VL-7B Bai et al. (2025) 41.21 45.82 38.77 41.80
@@ 4.1 Dataset and Metrics
Dataset. We evaluate on the validation and test sets of EarthReason Li et al. (2025). Pixel-level tasks directly utilize the original annotations, while region-level and object-level tasks are transformed using M2B Liu et al. (2024a) and M2C Miao et al. (2025) annotation formats based on the original labels, respectively. Additionally, we employ the RSVG Zhan et al. (2023) test set for Visual Grounding evaluation and the RRSISD Liu et al. (2024b) test set for Referring Expression Segmentation tasks, with their training sets explicitly excluded from model training. For unseen category evaluation, we manually selected 100 images covering 10 distinct unseen categories from the RemoteSAM-270K Yao et al. (2025b) dataset.
Metrics. Following previous works Yao et al. (2025b), Li et al. (2025), we adopt $gIoU$, cIoUcIoU, and $Acc@0.5$ as region- and pixel-level evaluation metrics. For contour reasoning, we employ F1 score with 1-pixel and 3-pixel dilated kernel ($F1@1$ & $F1@3$) as evaluation metrics, and incorporate two geometric error measures: Average Symmetric Distance ($ASD$) and Hausdorff Distance ($HD$).
Table 3: Geospatial Contour Reasoning Results on EarthReason.
Method $F1@1\uparrow$ $F1@3\uparrow$ $ASD\downarrow$ $HD\downarrow$
Val Test Val Test Val Test Val Test
EPOC Chen et al. (2024a) 0.10 0.10 0.19 0.20 60.07 64.38 243.98 255.81
RemoteReasoner 0.46 0.44 0.55 0.58 41.50 45.20 160.56 165.11

### v3

@@ Abstract
Remote sensing imagery presents vast, inherently unstructured spatial data, necessitating sophisticated reasoning to interpret complex user intents and contextual relationships beyond simple recognition tasks. In this paper, we aim to construct an Earth observation workflow to handle complex queries by reasoning about spatial context and user intent. As a reasoning workflow, it should autonomously explore and construct its own inference paths, rather than being confined to predefined ground‑truth sequences. Ideally, its architecture ought to be unified yet generalized, possessing capabilities to perform diverse reasoning tasks through one model without requiring additional fine-tuning. Existing remote sensing approaches rely on supervised fine-tuning paradigms and task‑specific heads, limiting both autonomous reasoning and unified generalization. To this end, we propose RemoteReasoner, a unified workflow for geospatial reasoning. The design of RemoteReasoner integrates a multi-modal large language model (MLLM) for interpreting user instructions and localizing targets, together with task transformation strategies that enable multi-granularity tasks, including object-, region-, and pixel-level. In contrast to existing methods, our framework is trained with reinforcement learning (RL) to endow the MLLM sufficient reasoning autonomy. At the inference stage, our transformation strategies enable diverse task output formats without requiring task-specific decoders or further fine-tuning. Experiments demonstrated that RemoteReasoner achieves state-of-the-art performance across multi-granularity reasoning tasks. Furthermore, it retains the MLLM’s inherent generalization capability, demonstrating robust performance on unseen tasks and categories.
Codehttps://github.com/1e12Leon/RemoteReasoner
@@ 1 Introduction
Remote sensing analysis has evolved beyond simple object detection (liu2025boost; li2025rsvg) or classification (adegun2023review), increasingly requiring stronger capabilities to support complex real-world decision-making. Although existing approaches (yao2025remotesam; liu2024remoteclip) are effective for multiple tasks, they struggle to interpret nuanced user intent and infer implicit spatial relationships for intricate queries. Reasoning (jaech2024openai; xu2025towards) bridges this gap by enabling systems to dynamically decipher the underlying goals of ambiguous or high-level instructions (e.g., “Find areas at risk of flooding near critical infrastructure after heavy rain”). Such a geospatial workflow would conveniently accommodate varied geoscience applications (zhao2024artificial).
@@
Figure 1: Comparison between Existing Frameworks and RemoteReasoner. Existing remote sensing reasoning approach (e.g., SegEarth-R1) requires SFT with annotated reasoning processes, and is limited to single-task outputs and task decoder. In contrast, our framework supports unsupervised reasoning and multi-granularity tasks.
@@
Numerous efforts (lai2024lisa; ren2024pixellm; chen2024sam4mllm) have explored reasoning tasks in the natural image domain. For instance, LISA (lai2024lisa) utilizes multi-modal large language models (MLLMs) (carolan2024review) for reasoning-driven segmentation tasks. However, these approaches rely on supervised fine-tuning (SFT) paradigms to train the models. Such frameworks constrain model autonomy, ultimately compromising generalization capabilities. This situation can be manifested in substantial performance degradation on out-of-distribution samples and catastrophic forgetting of general capabilities (liu2025seg). Similarly, the remote sensing community has seen pioneering works (li2025segearth; zhang2023spatial; wang2024earthvqa). SegEarth-R1 (li2025segearth) introduces a geospatial pixel reasoning task that integrates a hierarchical MLLM and a mask decoder to generate both reasoning processes and segmentation outputs. However, it remains constrained by SFT paradigms and exclusively supports segmentation tasks.
@@
These limitations motivate us to design a reasoning workflow that autonomously interprets user intent through self-directed thinking. It should also efficiently generate geospatial interpretations in diverse output formats, which are tailored for various downstream intelligence analysis. Firstly, the requirement for self-thinking should be fulfilled through an open-ended learning paradigm. This aligns naturally with the emerging reinforcement learning (RL) frameworks (guo2025deepseek; shen2025vlm). Rather than SFT approaches that demand strictly annotated reasoning processes, RL explores the model’s self-evolution potential to cultivate reasoning capabilities. Secondly, to flexibly support multi-granularity tasks, we should avoid redundant task decoders and training through decision-level task transformation.
@@
Followed by these insights, we propose RemoteReasoner, a reasoning workflow that integrates a robust MLLM and a flexible inference strategy. As shown in Fig. 1, building upon the pixel-level reasoning task proposed in SegEarth-R1 (li2025segearth), RemoteReasoner further implements two novel tasks through autonomous reasoning: 1) region-level: geospatial region reasoning, and 2) object-level: geospatial contour reasoning. These three distinct tasks can be achieved through a single MLLM inference followed by a unified decision-level task transformation pipeline, significantly improving output efficiency for multi-granularity reasoning tasks.
@@
During training, we utilize Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) to fine-tune the reasoning model. Benefiting from our sophisticated reward function that balances output formats, localization accuracy, and quantity, RemoteReasoner demonstrates strong capabilities in autonomous reasoning and precise region-level localization. This training approach maintains the inherent generalization capability of MLLM, thereby endowing it with the ability to recognize out-of-distribution categories. During inference, we design a task transformation pipeline inspired by RemoteSAM (yao2025remotesam), which enables the model to perform object- and pixel-level reasoning tasks, eliminating redundant computations across granularities.
@@
Extensive experimental results demonstrate that RemoteReasoner achieves state-of-the-art performance on multi-granularity reasoning tasks, evidencing its enhanced autonomous reasoning capabilities and superior generalization. For instance, RemoteReasoner achieves 3.67% accuracy improvement over SegEarth-R1 in pixel-level geospatial reasoning tasks. On unseen tasks (e.g., Referring Segmentation), our framework significantly outperforms competing models (e.g., PixelLM, LISA) in performance.
@@ 2.1 Reasoning in Large Language Models
Recent efforts (jaech2024openai; guo2025deepseek; ren2024pixellm; lai2024lisa; li2025segearth) have shown that reasoning is central to aligning large language models (LLMs) with complex instructions. Models like OpenAI-o1 (jaech2024openai) and DeepSeek-R1 (guo2025deepseek) incorporate chain-of-thought(CoT) (wei2022chain) reasoning into their generation process, enabling flexible and interpretable multi-step inference. Both adopt reinforcement learning (RL) (kaelbling1996reinforcement; li2017deep; arulkumaran2017deep) to induce generalizable reasoning behaviors—OpenAI-o1 through RLHF and CoT monitoring, and DeepSeek-R1 via large-scale RL from scratch followed by multi-stage pretraining. These approaches exhibit strong reasoning capabilities but remain limited to natural language and vision tasks.
In remote sensing, models such as PixelLM (ren2024pixellm), LISA (lai2024lisa), and SegEarth-R1 (li2025segearth) extend reasoning to spatial domains. While effective for pixel- or region-level inference, they often support narrow task types and fixed granularity. Notably, SegEarth-R1 relies on supervised fine-tuning, making it less robust to distribution shifts and less adaptable in open-ended settings. These limitations underscore the need for a more versatile geospatial framework.
Figure 2: Overview of our RemoteReasoner. We utilize GRPO (shao2024deepseekmath) to explore the model’s self-thinking capability. Then we design an inference workflow to perform multi-granularity reasoning tasks.
@@ 2.2 Remote Sensing Multi-modal Models
Recent remote sensing multimodal models (liu2024remoteclip; zhang2024rs5m; hu2025rsgpt; ma2025unified; li2024show; li2024languag) demonstrate strong capabilities in handling diverse tasks across spatial granularities. Falcon (yao2025falcon) supports 14 vision-language tasks via multi-resolution encoding; GeoChat (kuckreja2024geochat) and LHRS-Bot (muhtar2024lhrs) incorporate spatial prompts and geographic priors for improved region-level grounding; RemoteSAM (yao2025remotesam) adopts a unified encoder for dense pixel-level segmentation; SkyEyeGPT (zhan2024skyeyegptunifyingremotesensing) reformulates task input/output as natural language for broad task generalization. While effective in geospatial understanding, these models lack robust reasoning capabilities, limiting performance under ambiguous or compositional prompts. While existing models demonstrate strong perception capabilities, they lack the reasoning ability to interpret complex prompts or adapt across granularities.
@@ 3.1 Data
The EarthReason (li2025segearth) dataset is proposed alongside SegEarth-R1, containing 6 questions per sample with an average length of 20.86 words. It comprises 2,371 training, 1,135 validation, and 1,928 testing samples. However, this dataset is limited to pixel tasks. To address this, we generate detection and contour annotations through Mask2Bbox (M2B) (liu2024remoteclip) and Mask2Contour (M2C) (miao2025prompting) strategies, respectively. Additionally, we design prompt templates combined with original questions and generated bounding boxes to form training samples. The segmentation and contour annotations are exclusively reserved for downstream testing evaluation.
@@ 3.2 Reasoning Model
We utilize Qwen2.5-VL-7B-instruct (bai2025qwen2) as our reasoning model $\mathcal{F}_{\theta}$. Despite Qwen2.5-VL’s superior performance in general-domain applications, direct adaptation to remote sensing imagery introduces domain knowledge discrepancy. Moreover, its reasoning capabilities require significant enhancement to address complex geospatial tasks. Therefore, we require fine-tuning it through reinforcement learning on remote sensing data.
@@ 3.3 Training
We utilize Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) to train RemoteReasoner’s reasoning model. Unlike traditional reinforcement learning methods like PPO (schulman2017proximal), GRPO directly compares groups of candidate outputs without an additional critic model. Given a query $Q$, GRPO samples N candidate outputs $o_{1},o_{2},...,o_{N}$ from the policy $\pi_{\theta}$. Then it computes each output $o_{i}$ through a Reward function $R(Q,o_{i})$. To determine the relative quality of the outputs, GRPO further normalizes the obtained rewards and subsequently derives the advantage:
@@
$\frac{1}{N}\sum_{i=1}^{N}\{min[c_{1}\cdot A_{i},c_{2}\cdot A_{i}]$
$-\beta\mathbb{D}_{KL}[\pi_{\theta}||\pi_{ref}]\},$
@@ 3.4 Inference Workflow
After training, the reasoning model itself generates bounding boxes (bbox) indicating target regions. To further enhance the workflow and enable both pixel-level and object-level reasoning tasks, we need to leverage established tools. Specifically, we first input the bbox and its centroid coordinates to the SAM2 (ravi2024sam) to generate pixel-level masks. Subsequently, morphological operations are applied to analyze the mask images, extracting the outermost polygonal vertices to obtain target contours. The mathematical formulation of this process is presented as follows:
@@ 4 Experiments
LISA 57.39 59.10 61.04 60.88
PixelLM 57.79 59.22 57.94 60.01
PSALM 62.03 64.61 66.61 68.30
SegEarth-R1 64.13 68.25 68.60 70.75
@@
DeepSeek-VL2-tiny 12.08 12.67 17.51 18.62
GeoChat 10.10 8.89 12.57 11.44
Qwen2.5-VL-7B 41.21 45.82 38.77 41.80
@@
Table 2: Geospatial Region Reasoning Results.
Figure 3: Geospatial Contour Reasoning Results. We select EPOC (chen2024subobject) for comparison.
Method $F1@1\uparrow$ $F1@3\uparrow$ $Recall@1\uparrow$ $Recall@3\uparrow$ $ASD\downarrow$ $HD\downarrow$
Val Test Val Test Val Test Val Test Val Test Val Test
EPOC 0.10 0.10 0.19 0.20 0.14 0.15 0.28 0.29 60.07 64.38 243.98 255.81
RemoteReasoner 0.46 0.44 0.55 0.58 0.40 0.47 0.53 0.64 41.50 45.20 160.56 165.11
Table 3: Geospatial Contour Reasoning Results on EarthReason.
@@ 4.1 Dataset and Metrics
Dataset. We evaluate on the validation and test sets of EarthReason (li2025segearth). Pixel-level tasks directly utilize the original annotations, while region-level and object-level tasks are transformed using M2B (liu2024remoteclip) and M2C (miao2025prompting) annotation formats based on the original labels, respectively. Additionally, we employ the RSVG (zhan2023rsvg) test set for Visual Grounding evaluation and the RRSISD (liu2024rotated) test set for Referring Expression Segmentation tasks, with their training sets explicitly excluded from model training. For unseen category evaluation, we manually selected 100 images covering 10 distinct unseen categories from the RemoteSAM-270K (yao2025remotesam) dataset.
Metrics. Following previous works (yao2025remotesam; li2025segearth), we adopt $gIoU$, cIoUcIoU, and $Acc@0.5$ as region- and pixel-level evaluation metrics. For contour reasoning, we followed (jiao2025roipoly) to employ F1 score with 1-pixel and 3-pixel dilated kernel ($F1@1$ & $F1@3$) as evaluation metrics, and incorporate two geometric error measures: Average Symmetric Distance ($ASD$) and Hausdorff Distance ($HD$).